Allison:

Rewrite the expression without using grouping symbols. 12(x-17)

Pixel:

Do they want it in word form?

Allison:

12x-17??

Allison:

Nope~

Allison:

12x-17 12x-12*17 12+12*17 12x-29

Allison:

@Shadow

Shadow:

I don't know how to do this

Shadow:

Distributive Property A number multiplied by a parentheses with term(s) inside will be multplied to each term within the parentheses.

Shadow:

What do you think the answer is?

Shadow:

\[12x - 204 = (12 \times x) - (12 \times 17)\]

Allison:

So B

Shadow:

Correct

Allison:

13x-29x 13(x+29) 13(x-29) x(13-29) x(29-13)

Shadow:

Hmm, what do you think about this one?

Allison:

A?

Shadow:

\[13(x + 29) = (13 \times x) + (13 \times 29)\]

Allison:

So yes

Allison:

Oh nvm is it B

Shadow:

Let me illustrate a concept to you. \[(8x + 20) = 4(2x + 5) \] You can pull numbers out of a parentheses to simplify

Shadow:

Don't guess

Allison:

I have 7 minutes left to do this, ain't nobody got time for that. D:

Shadow:

\[13(x-29) = (13 \times x) - (13 \times 29)\]

Allison:

And we're barely on the second question.

Shadow:

Look at this: \[(3h + 7h) = h(3 + 7)\]

Allison:

It is B

Shadow:

It is not

Allison:

C

Shadow:

Yes, but I do not think you understand the rule ._.

Allison:

I break rules. You're a Mod, you should know I do xD

Shadow:

Not thee rule, I allude to the Distributive Property

Allison:

Eh.

Allison:

Z/3-4z+4(z-5) when z=15 45?

Shadow:

\[\frac{ 15 }{ 3 } -4(15) + 4(15 -5)\] \[5 - 60 + 40 = -15\]

Allison:

Oh ._.

Allison:

Which shows how to use the distributive property to evaluate 6*72? 6*(70+8) 6*(70+2) 6*(70-2) 6*(70-8)

Shadow:

Which one makes the MOST sense?

Allison:

Lmao. B.

Shadow:

Yes ._.

Allison:

Kms.

Allison:

Identify the like terms. 3q, -3p, 3pq, -11pq 3q,-3p, and 3pq -3p and 3pq 3 q and 3 pq 3pq and -11pq

Shadow:

Like terms means that they can be added and subtracted by each other. You cannot add terms that have a differing variable to the first. Example: 3x - 5y This cannot occur. But, 3xy - 5xy can

Shadow:

You would get -2xy

Allison:

A?

Shadow:

3q - 3p + 3pq cannot occur. These are not like terms. They do not have the same variables.

Allison:

C

Shadow:

q and pq are not the same

Allison:

D

Shadow:

pq and pq are the same

Allison:

So yes

Shadow:

Yes ._.

Allison:

2/3a + 4/3b - 4/3a 2/3a 4/3b - 8/9a 2/3ab 4/3b - 2/3 a

Shadow:

This problem wants us to illustrate our knowledge of like terms. Identify for me what terms we can add

Allison:

Combine like terms e.e

Allison:

I'm gonna guess on it

Shadow:

Why ._.

Allison:

Because I'm timed.

Shadow:

Terms with a can only add/subtract with terms with a Terms with b can only add/subtract with terms with b

Shadow:

\[\frac{ 2 }{ 3 }a + \frac{ 4 }{ 3 }b - \frac{ 4 }{ 3}a\] Let me rearrange it for you \[\frac{ 2 }{ 3 }a - \frac{ 4 }{ 3}a + \frac{ 4 }{ 3 }b \]

Shadow:

You do 2/3a - 4/3a, then leave b alone since you cannot do anything with it(no other like terms, terms with b in it)

Allison:

D?

Shadow:

Yes